# Efficient Computation of Whole-Body Control Utilizing Simplified Whole-Body Dynamics via Centroidal Dynamics

### v1

@@ Abstract
I INTRODUCTION
The field of humanoid robotics has witnessed significant advances in recent years, driven by the demand for robots that can operate in environments designed for humans and perform tasks ranging from simple repetitive actions to complex, dynamic interactions. One of the essential components in realizing it is the development of Whole-Body Control (WBC) frameworks, as initially explored in foundational research [1], [2]. WBC facilitates the execution of multi-task control at the torque level within the operational space, carefully accounting for contact dynamics and ensuring consistency with various constraints. Following its initial implementation [3], research efforts have continued to advance the control of humanoid robots with the implementation of the WBC framework using experiment [4], [5], [6].
@@ I INTRODUCTION
To elucidate these contributions, our paper is structured as follows: In Section II, a brief overview of the whole-body control knowledge is provided. Section III proposes new novel method of constructing reduced dynamics. In Section IV the application of proposed method to conventional methods are introduced. The section V provides the result with the proposed method. Finally, Section VI concludes the paper with a discussion based on the validation findings.
II Constraint for Whole-Body Control
@@
$\mathbf{M}(\mathbf{q})\mathbf{\ddot{q}}+\mathbf{b}(\mathbf{q},\mathbf{\dot{q}})+\mathbf{g}(\mathbf{q})+\mathbf{J}_{c}(\mathbf{q})^{T}\mathbf{F}_{c}=\mathbf{S}^{T}\boldsymbol{\Gamma},$ (1)
where $\mathbf{q}\in\mathbb{R}^{n+6}$ is the generalized joint position vector of the system, and $n$ is the degrees of freedom (DOF) of the robot actuated joints. $\mathbf{M}(\mathbf{q})\in\mathbb{R}^{(n+6)\times(n+6)}$ is the inertia matrix, $\mathbf{b}(\mathbf{q},\mathbf{\dot{q}})\in\mathbb{R}^{n+6}$ is the Coriolis and centrifugal force vector, $\mathbf{g}(\mathbf{q})\in\mathbb{R}^{n+6}$ is the gravity force vector, $\mathbf{S}^{T}\in\mathbb{R}^{{(n+6)}\times n}$ is a selection matrix that specifies actuated joints from system joints, $\mathbf{J}_{c}(\mathbf{q})\in\mathbb{R}^{c\times(n+6)}$ is the constraint Jacobian, which usually used to describe the contact constraint of the robot. $c$ is the constrained DOF, and $\mathbf{F}_{c}\in\mathbb{R}^{c}$ is a vector that concatenate all constriant forces.
The iWBC solves the optimization problem that considers whole-body dynamics to generate a feasible joint control solution when task space trajectories are given. Various inequality and equality constraints must be considered for the iWBC optimization problem. For a humanoid robot interacting with the external environment, the necessary constraints include the equality constraint with (1) to describe the robot’s system dynamics. Additionally, the following inequality and equality constraints are required:
@@
$\mathbf{M}_{l}\mathbf{\ddot{q}}+\mathbf{b}_{l}+\mathbf{g}_{l}+\mathbf{J}_{c,l}^{T}\mathbf{F}_{c}$ $=\boldsymbol{\Gamma}$ (7)
@@
Equation (6) represents the top six equations of (1), and Equation (7) comprises the bottom $n$ equations of (1). Equation (6) describes the Newton-Euler equations for the floating base and can be used as the dynamics constraint instead of (1). Through Equation (7), control torques, $\boldsymbol{\Gamma}$ can be calculated with $\ddot{\mathbf{q}}$ and $\mathbf{F}_{c}$. Therefore, by excluding torque from the decision variable $\mathbf{x}$ the optimization problem can be solved efficiently.
III CONSTRUCTION OF UNCONSTRAINED IMPLICIT DYNAMICS
III-A Categorization and Simplification of Kinematic Chains
@@
In our method, we categorize the kinematic chain of floating-base robots into three types, as shown in Fig. 1. The first type is the virtual chain, which consists of virtual joints that describe the six DOF floating base body motion relative to the inertial frame. The second type is the $n_{cc}$-DOF constraint chain, which includes the kinematic chain between the base link and the links in constraint. These constraint links are typically described by $\mathbf{J}_{c}$, forming the constrained space (usually used as the contact space for walking robot). The third type is the $n_{uc}$-DOF unconstrained chain, which is the kinematic chain composed of the robot’s links excluding the links included in the constraint chain.
In the proposed framework, the dynamical model of the unconstrained chain is simplified into six DOF utilizing the centroidal dynamics concept. Then, the $n$-DOF robot model can be reduced to $(n_{cc}+12)$-DOFs. The number 12 is from the summation of six DOF simplified unconstrained chain and six DOF virtual chain. The relationship between the original robot model and the reduced model can be seen in Fig. 2. Then the unconstrained chain is considered sequentially, resulting in the whole dynamics ($n+6$)-DOF being divided into two sequential problems: the reduced dynamics $(n_{r}+6)$-DOF and the unconstrained chain dynamics $(n_{uc})$-DOF. The $n_{r}$ denotes the DOF of actuated joints of reduced dynamics, which is $n_{cc}+6$. Therefore, the proposed method decreases whole-body controller computational cost by reducing the DOF of the system dynamics and decision variable size.
Figure 1: Schematic representation of the robot model system, depicting the virtual chain, the constraint chain, and the unconstrained chain. In case of (a), the lower body forms the constraint chain while the upper body forms the unconstrained chain. In (b) the connected rigid bodies from the foot to the hand constitute the constraint chain, while the remaining parts—head, the other arm, and leg—form the unconstrained chain.
@@
Figure 2: Composition comparison between the ordinary robot model and the reduced-dimension robot model. The gray box on the left denotes the rigid-body dynamics model, and the gray box on the right denotes the reduced dimensional robot model. $n$ is the DOF of the actuated joints of the original robot model, $n_{r}$ is the DOF of the actuated joints of reduced dimensional model. $n_{vc}$, $n_{cc}$, $n_{uc}$ represent the DOF of the virtual chain, constraint chain, and unconstrained chain, respectively. $n_{s}$ is the DOF of the simplified unconstrained chain, which is six.
III-B Interpreting Unconstrained Chain with Centroidal dynamics
@@
$\mathbf{h}_{G,uc}=\mathbf{M}_{G,uc}(\mathbf{q})\mathbf{\dot{q}},$ (8)
@@
Where $\mathbf{I}_{G,uc}\in\mathbb{R}^{6\times 6}$ is the inertia matrix of the unconstrained chain.
@@
III-C Construction of Reduced Whole-Body Dynamics
@@
$\mathbf{\dot{q}}_{r}={\begin{bmatrix}{\mathbf{\dot{q}}_{vc}}\\ {\mathbf{\dot{q}}_{cc}}\\ {{{}^{b}}\mathbf{v}_{G,uc}}\end{bmatrix}},$ (11)
where $\mathbf{\dot{q}}_{vc}\in\mathbb{R}^{6}$ is the generalized velocity of the virtual chain joints from an inertial frame, $\mathbf{\dot{q}}_{cc}\in\mathbb{R}^{n_{co}}$ is the generalized joint velocity of the constraint chain joints. The superscript preceding a variable denotes its reference frame. Here, ${}^{b}\mathbf{v}_{G,uc}$ indicates that $\mathbf{v}_{G,uc}$ is the spatial velocity of unconstrained chain relative to the reference frame of the robot base.
@@
$\mathbf{J}_{r}=\begin{bmatrix}\mathbf{S}_{vc}\\ \mathbf{S}_{cc}\\ {{}^{b}}\mathbf{J}_{G,uc}\mathbf{S}_{uc}\end{bmatrix},$ (13)
@@
$\mathbf{M}_{r}\mathbf{\ddot{q}}_{r}+\mathbf{b}_{r}+\mathbf{g}_{r}+\mathbf{J}_{c,r}^{T}\mathbf{F}_{c}=\mathbf{S}_{r}^{T}\boldsymbol{\Gamma}_{r},$ (14)
@@
$\mathbf{J}_{c,r}^{T}=\begin{bmatrix}\mathbf{S}_{vc}\\ \mathbf{S}_{cc}\\ \mathbf{0}_{6\times(n+6)}\end{bmatrix}\mathbf{J}_{c}^{T},$ (15)
instead of (14e) since the contact Jacobian matrix inherently does not include components from the unconstrained chain.
Consequently, (14) can be considered a simplified representation of a system interacting with the external environment, and it can be utilized as equations of motion constraint for the reduced model in the iWBC method.
The effect of the unconstrained chain on the external environment can be described using (14), following the simplification process. Since the simplified model cannot capture the internal motion of the unconstrained chain, it’s dynamics must also be described. By considering the unconstrained chain as an independent model, it’s dynamics are given by:
$\mathbf{M}_{uc}\ddot{\mathbf{q}}_{uc}+\mathbf{b}_{uc}+\mathbf{g}_{uc}=\boldsymbol{\Gamma}_{uc},$ (16)
@@
III-D Motion Control Tasks via Reduced Dynamics Model
@@
III-D1 Task Constraints in the Reduced Model Joint Space
@@
III-D2 Task Constraints in the Unconstrained Chain
@@
III-D3 Reference Acceleration Constraint
@@
IV Problem Formulation
IV-A Lexicographic Quadratic Programming
@@
where $\mathbf{x}=[\,\mathbf{\ddot{q}}^{T}\>\mathbf{F}_{c}^{T}\>\boldsymbol{\Gamma}^{T}]^{T}$, $\mathbf{A}\in\mathbb{R}^{m_{a}\times(2n+6+6c)}$, $\mathbf{a}\in\mathbb{R}^{m_{a}}$, $\mathbf{B}\in\mathbb{R}^{m_{b}\times(2n+6+6c)}$, $\mathbf{b}\in\mathbb{R}^{m_{b}}$. $m_{a}$ and $m_{b}$ are the total number of inequality and equality constraints of all priorities, respectively. $c$ is the number of contact links.
@@
IV-B LQP Formulation with Reduced Dynamics
@@
Equations (27a) and (27b) represent the constraints mentioned in Section II, formulated through reduced dynamics. For the task space, the constraints are constructed using the reduced dynamics, as discussed in subsection III-D, which can only be composed through such task constraints.
@@
Figure 3: The algorithm flow chart utilizing the proposed reduced dynamics.
@@
The constraint chain control torque $\boldsymbol{\Gamma}_{cc}^{*}\in\mathbb{R}^{n_{cc}}$ from the computation result of LQP 1 and the unconstrained chain control torque $\boldsymbol{\Gamma}_{uc}^{*}\in\mathbb{R}^{n_{uc}}$ from the LQP 2 are used to construct the control torque input of the robot $\boldsymbol{\Gamma}^{*}=[{\boldsymbol{\Gamma}_{cc}^{*}}^{T}~{}{\boldsymbol{\Gamma}_{uc}^{*}}^{T}]^{T}$.
V Evaluation
V-A System Setup
To validate the efficiency and performance of the method proposed in this paper, we employed the two-level computation framework to configure an HQP controller and compared its performance with that of a conventional HQP controller. The proposed framework was validated through the simulation. The robot used in the simulation, TOCABI, is human-sized (1.8m in height, 100kg in weight), 33-DOF torque-controlled humanoid robot. It is equipped with eight DOFs in each arm, six in each leg, three in the waist, and two in the head. The experiment was conducted on i7-10700 CPU, with the simulator MuJoco. For the QP solver, one of the off-the-shelf solvers, OSQP, was used. The robot’s operating system is Ubuntu 20.04, with a real-time patch applied using Xenomai 3.2.1. Further detailed information about the robot and system configuration can be found in [38].
TABLE I: Task Configuration with Conventional Method

### v2

@@ Abstract
keywords:
Dynamics, humanoid robot, model simplification, whole-body control.
[…]
Manuscript received January 10, 2021; revised March 10, 2021; accepted May 10, 2021. Recommended by Associate Editor Soon-Shin Lee under the direction of Editor Milton John. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2021R1A2C3005914). It was also supported by TUM AGENDA 2030, funded by the Federal Ministry of Education and Research (BMBF) and the Free State of Bavaria under the Excellence Strategy of the Federal Government and the Länder as well as by the Hightech Agenda Bavaria.
J. Ahn, H. Lee and J. Park are with the Graduate School of Convergence Science and Technology, Seoul National University (SNU), Gwanak-ro 1, Gwanak-gu, Seoul, Republic of Korea. (e-mails: {june992, hkleetony, park73}@snu.ac.kr). J. Jung is with Global Infrastructure and manufacturing, Samsung Electronics. (e-mails: jays.jung@samsung.com). Y. Lee is with the Korea Institute of Science and Technology (KIST), Seoul, 02792, Republic of Korea (e-mails: yisoo.lee@kist.re.kr). S. Haddadin is with Technical University of Munich, Germany; TUM School of Computation, Information and Technology (CIT); Chair of Robotics and Systems Intelligence (RSI); Munich Institute of Robotics and Machine Intelligence (MIRMI) (e-mails: haddadin@tum.de). J. Park is also with ASRI, RICS, Seoul National University, Republic of Korea, and Advanced Institutes and Advanced Institutes of Convergence Technology(AICT).
* Corresponding author.
Junewhee Ahn, Jaesug Jung, Yisoo Lee, Hokyun Lee, Sami Haddadin and Jaeheung ParkManuscript Template for the International Journal of Control, Automation, and Systems: ICROS & KIEExxxxxxxx
1 INTRODUCTION
The field of humanoid robotics has witnessed significant advances in recent years, driven by the demand for robots that can operate in environments designed for humans and perform tasks ranging from simple repetitive actions to complex, dynamic interactions. One of the essential components in realizing this is the development of Whole-Body Control (WBC) frameworks, initially explored in foundational research [1, 2]. WBC facilitates the execution of multi-task control at the torque level within the operational space, carefully accounting for contact dynamics and ensuring consistency with various constraints.
@@ 1 INTRODUCTION
To elucidate these contributions, the paper is structured as follows: Section 2 provides an overview of whole-body control. Section 3 introduces a method for constructing reduced dynamics. Section 4 discusses the application of the proposed method to conventional approaches. Section 5 presents the results. Finally, Section 6 concludes the paper with a discussion based on the validation findings.
2 CONSTRAINT FOR WHOLE-BODY CONTROL
@@
$\mathbf{M}(\mathbf{q})\mathbf{\ddot{q}}+\mathbf{b}(\mathbf{q},\mathbf{\dot{q}})+\mathbf{g}(\mathbf{q})+\mathbf{J}_{c}(\mathbf{q})^{T}\mathbf{F}_{c}=\mathbf{S}^{T}\bm{\Gamma},$ (1)
where $\mathbf{q}\in\mathbb{R}^{n+6}$ is the generalized joint position vector of the system, and $n$ is the degrees of freedom (DOF) of the robot actuated joints. $\mathbf{M}(\mathbf{q})\in\mathbb{R}^{(n+6)\times(n+6)}$ is the inertia matrix, $\mathbf{b}(\mathbf{q},\mathbf{\dot{q}})\in\mathbb{R}^{n+6}$ is the Coriolis and centrifugal force vector, $\mathbf{g}(\mathbf{q})\in\mathbb{R}^{n+6}$ is the gravity force vector, $\mathbf{S}^{T}\in\mathbb{R}^{{(n+6)}\times n}$ is a selection matrix that specifies actuated joints from system joints, $\mathbf{J}_{c}(\mathbf{q})\in\mathbb{R}^{c\times(n+6)}$ is the constraint Jacobian, which usually used to describe the contact constraint of the robot. $c$ is the constrained DOF, and $\mathbf{F}_{c}\in\mathbb{R}^{c}$ is a vector that concatenate all constraint forces. The iWBC solves the optimization problem that considers whole-body dynamics to generate a feasible joint control solution when task space trajectories are given. Various inequality and equality constraints must be considered for the iWBC optimization problem. For a humanoid robot interacting with the external environment, the necessary constraints include the equality constraint with (1) to describe the robot’s system dynamics. Additionally, the following inequality and equality constraints are required:
@@
$\mathbf{M}_{l}\mathbf{\ddot{q}}+\mathbf{b}_{l}+\mathbf{g}_{l}+\mathbf{J}_{c,l}^{T}\mathbf{F}_{c}$ $=\bm{\Gamma}$ (7)
@@
Equation (6) represents the top six equations of (1), and Equation (7) comprises the bottom $n$ equations of (1). Equation (6) describes the Newton-Euler equations for the floating base and can be used as the dynamics constraint instead of (1). Through Equation (7), control torques, $\bm{\Gamma}$ can be calculated with $\ddot{\mathbf{q}}$ and $\mathbf{F}_{c}$. Therefore, by excluding torque from the decision variable $\mathbf{x}$ the optimization problem can be solved efficiently.
3 CONSTRUCTION OF UNCONSTRAINED IMPLICIT DYNAMICS
3.1 Categorization and Simplification of Kinematic Chains
@@
In the proposed method, the kinematic chain of floating-base robots is divided into three types. The virtual chain consists of virtual joints that represent the robot’s floating base relative to the inertial frame. The constraint chain includes the links from the floating base to the constraint link. The constraint links are typically described by $\mathbf{J}_{c}$, forming the constrained space (usually used as the contact space for walking robot). The unconstrained chain comprises the links not included in the constraint chain. Additionally, the constraint chain and the unconstrained chain are composed of joints required to describe the motion of each link from the frame of the floating base. The DOF of these joints are $n_{cc}$ and $n_{uc}$, respectively.
Fig. 1 illustrates this classification of kinematic chains using the simplified representation of robot. In Fig. 1 (a), the kinematic chain of the constraint chain and the unconstrained chain are clearly separated from the floating base, making the distinction straight forward. However, in Fig. 1 (b), the left arm and neck forming the unconstrained chain are not directly connected to the floating base. In this case, the waist joint, which is included in the constraint chain, is also required for the unconstrained chain to describe the motion from the frame fo the floating base.
In the proposed framework, the dynamical model of the unconstrained chain with more than six DOF is simplified into six DOF utilizing the centroidal dynamics. The relationship between the original robot model and the reduced model can be seen in Fig. 2. By reducing the $n_{uc}$ DOF of the unconstrained chain to 6 DOF, the robot’s model DOF can be simplified from $n$ to $n_{cc}+6$. This reduction significantly decreases computational complexity in control algorithms, especially for robots with a large number of joints.
While the kinematic chain is divided into two types based on constraints in this paper, the proposed method is not limited to this constraint specific criterion. The crucial point is that it compresses the DOF of one of the chains. Therefore, the designation of the floating base or the scope of the kinematic chain to be compressed can be adjusted according to the user’s goals to enhance computational efficiency.
Figure 1: Simplified representation of the robot’s links, joints, illustrating the categorization into virtual chain, constraint chain, and unconstrained chain.
@@
Figure 2: Composition comparison between the full model and the reduced-dimension model. $n_{vc}$, $n_{cc}$, $n_{uc}$ represent the DOF of the virtual chain, constraint chain, and unconstrained chain, respectively.
3.2 Interpreting Unconstrained Chain with Centroidal dynamics
@@
$\mathbf{h}_{G,uc}=\mathbf{M}_{G,uc}(\mathbf{q}_{uc})\dot{\mathbf{q}}_{uc},$ (8)
@@
Where $\mathbf{I}_{G,uc}\in\mathbb{R}^{6\times 6}$ is the centroidal inertial matrix of the simplified unconstrained chain.
@@
3.3 Construction of Reduced Whole-Body Dynamics
@@
$\mathbf{\dot{q}}_{r}={\begin{bmatrix}{\mathbf{\dot{q}}_{vc}^{T}}&{\mathbf{\dot{q}}_{cc}^{T}}&{{{}^{b}}\mathbf{v}_{G,uc}^{T}}\end{bmatrix}}^{T},$ (11)
where $\mathbf{\dot{q}}_{vc}\in\mathbb{R}^{6}$ is the generalized velocity of the virtual chain joints from an inertial frame, $\mathbf{\dot{q}}_{cc}\in\mathbb{R}^{n_{cc}}$ is the generalized joint velocity of the constraint chain joints. The superscript preceding a variable denotes its reference frame. Here, ${}^{b}\mathbf{v}_{G,uc}$ indicates that $\mathbf{v}_{G,uc}$ is the spatial velocity of unconstrained chain relative to the reference frame of the robot base.
@@
$\mathbf{J}_{r}=\begin{bmatrix}\mathbf{S}_{vc}^{T}&\mathbf{S}_{cc}^{T}&({{}^{b}}\mathbf{J}_{G,uc}\mathbf{S}_{uc})^{T}\end{bmatrix}^{T},$ (13)
@@
$\mathbf{M}_{r}\mathbf{\ddot{q}}_{r}+\mathbf{b}_{r}+\mathbf{g}_{r}+\mathbf{J}_{c,r}^{T}\mathbf{F}_{c}=\mathbf{S}_{r}^{T}\bm{\Gamma}_{r},$ (14)
@@
$\mathbf{J}_{c,r}^{T}=\begin{bmatrix}\mathbf{S}_{vc}^{T}&\mathbf{S}_{cc}^{T}&(\mathbf{0}_{6\times(n+6)})^{T}\end{bmatrix}^{T}\mathbf{J}_{c}^{T},$ (15)
instead of (14e) since the contact Jacobian matrix inherently does not include components from the unconstrained chain. The vector $\bm{\Gamma}_{r}\in\mathbb{R}^{n_{r}}$ represents the actuated joint torques of the reduced model, which has the following relationship according to the OSF:
$\mathbf{S}^{T}\Gamma=\mathbf{J}_{r}^{T}\mathbf{S}_{r}^{T}\Gamma_{r}+\mathbf{N}_{r}^{T}\Gamma_{o},$ (16)
where $\mathbf{N}_{r}^{T}$ is the null-space projection matrix of $\mathbf{J}_{r}^{T}$ and $\mathbf{N}_{r}^{T}\Gamma_{o}$ represents the null-space torque vector that does not affect the reduced model space. Consequently, (14) can be considered a simplified representation of a system interacting with the external environment, and it can be utilized as equations of motion constraint for the reduced model in the iWBC method.
The effect of the unconstrained chain on the external environment can be described using (14), following the simplification process. Since the simplified model cannot capture the internal motion of the unconstrained chain, it’s dynamics must also be described. By considering the unconstrained chain as an independent model on the base link’s frame, it’s dynamics is same with the rigid body dynamics on the fixed frame;
$\mathbf{M}_{uc}\ddot{\mathbf{q}}_{uc}+\mathbf{b}_{uc}+\mathbf{g}_{uc}=\bm{\Gamma}_{uc},$ (17)
@@
3.4 Motion Control Tasks via Reduced Space
@@
3.4.1 Task Constraints in the Reduced Model Space
@@
3.4.2 Task Constraints in the Unconstrained Chain
@@
3.4.3 Reference Acceleration Constraint
@@
4 Problem Formulation
4.1 Lexicographic Quadratic Programming
@@
where $\mathbf{x}=[\,\mathbf{\ddot{q}}^{T}\>\mathbf{F}_{c}^{T}\>\bm{\Gamma}^{T}]^{T}$, $\mathbf{A}\in\mathbb{R}^{m_{a}\times(2n+6+6c)}$, $\mathbf{a}\in\mathbb{R}^{m_{a}}$, $\mathbf{B}\in\mathbb{R}^{m_{b}\times(2n+6+6c)}$, $\mathbf{b}\in\mathbb{R}^{m_{b}}$. $m_{a}$ and $m_{b}$ are the total number of inequality and equality constraints of all priorities, respectively. $c$ is the number of contact links.
@@
4.2 LQP Formulation with Reduced Dynamics
@@
Equations (28a) and (28b) represent the constraints mentioned in Section 2, formulated through reduced dynamics. For the task space, the constraints are constructed using the reduced dynamics, as discussed in subsection 3.4, which can only be composed through such task constraints.
@@
As mentioned in section 3.1, overlapping joints included in both chain can exist. In such cases, the acceleration of these joints obtained from the solution in LQP 1 are also considered as constraints in LQP 2.
Figure 3: The algorithm flow chart of the proposed method.
@@
The constraint chain control torque $\bm{\Gamma}_{cc}^{*}\in\mathbb{R}^{n_{cc}}$ from the computation result of LQP 1 and the unconstrained chain control torque $\bm{\Gamma}_{uc}^{*}\in\mathbb{R}^{n_{uc}}$ from the LQP 2 are used to construct the control torque input of the robot $\bm{\Gamma}^{*}=[{\bm{\Gamma}_{cc}^{*}}^{T}~{}{\bm{\Gamma}_{uc}^{*}}^{T}]^{T}$.
5 Evaluation
5.1 System Setup
The proposed method was validated through the simulation. The robot used in the simulation, TOCABI, is human-sized (1.8m in height, 100kg in weight), 33-DOF torque-controlled humanoid robot. It is equipped with eight DOFs in each arm, six in each leg, three in the waist, and two in the head. The experiment was conducted on i7-10700 CPU, with the simulator MuJoco. For the QP solver, one of the off-the-shelf solvers, OSQP, was used. The robot’s operating system is Ubuntu 20.04, with a real-time patch applied using Xenomai 3.2.1. Further detailed information about the robot and system configuration can be found in [34].
Table 1: Task Configuration with Conventional Method